OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 4(x - 6) = -2(x - 3)

OpenStudy (anonymous):

4(x-6)=-2(x-3) 4x - 24 = -2x + 6 Move variables to one side and numbers to one side! 4x + 2x = 24 + 6 6x = 30 x = 5

OpenStudy (anonymous):

thank you! :)

OpenStudy (anonymous):

Thanks!